# Rosette formations as symmetry-breaking events: theory and experiment

### v1

@@ Abstract
Multicellular rosettes are observed in different situations such as morphogenesis, wound healing, and cancer progression. While some molecular insights have been gained to explain the presence of these assemblies of five or more cells around a common center, what are the tunable, global features that favors/hinders their formation is still largely unknown. Here, we made use of a Voronoi dynamical model to investigate the ingredients driving the emergence of rosettes characterized by different degree of stability and organization. We found that (i) breaking the local spatial symmetry of the system, i.e., introducing curvature-inducing defects, allows for the formation of rosette-like structures (ii) whose probability of formation depends on the characteristic of the cellular layer. In particular, a trade-off between tissue fluidity and single cell deformability dictates the assembly of transient rosettes, that are strongly stabilized in the presence of cell alignment interactions.
To test our model predictions, we performed fluorescence microscopy experiments on rosette-forming neural populations derived from induced pluripotent stem cells (iPSCs), finding significant agreement. Overall, our work may set the stage to gain an unifying understanding of the plethora of biophysical mechanisms involving the occurrence of rosette-like structure both in physiology and their altered formation in pathology.
@@ I Introduction
Rosettes are ubiquitous structures in biological systems. These multicellular assemblies, composed of five or more cells that interface at a central point, are observed in stem cell populations in vitro and in vivo both during morphogenesis and in several adult-organism contexts Harding et al. (2014). In particular, polarized epithelial rosettes are evolutionary-preserved intermediate states for the organogenesis of multiple organs across species Harding et al. (2014). Despite their centrality, the environmental conditions, i.e. the extracellular signals and cellular events, driving rosette formation are yet to be fully determined Miotto et al. (2023). In addition, different kind of rosettes are observed Harding et al. (2014), which may further reflect in distinct ways in which they form. During development, two classes of rosettes are found with different structural features and formation patterns. The coordinated apical constriction of neighbor cells results in rosettes having cells with their apical domains pointed toward a common center. Another formation mechanism relies on the activity of planar polarized myosin, where planar polarization is obtained thanks to the asymmetric distribution of molecules along the cell axis Gibson and Perrimon (2003); Peruzzi et al. (2021); Caudo et al. (2024); Miotto et al. (2025).
@@
Figure 1: Modeling rosette formation. a) Examples of stationary configurations in the solid and fluid-like regimes are shown with cells colored in blue and red, respectively. b) (top) Probability density distribution (pdf) of the cells’ egdes lengths for the configurations shown in panel a). (bottom) Frequency of occurrence of cell as a function of their edges number. c) Sketches of three possible rosette conformations and the corresponding used model. (i) Transient rosette configuration, with 5 or more cells sharing a vertex are simulated breaking spatial symmetric via a harmonic potential acting on a given point. (ii) Rosette with a central lumen are obtained via a Lennard-Jones potential energy originating in the center of the lumen, while (iii) pinwheel rosette are simulated inserting in the simulation a leader cell having a different target area.
Siang and coworkers Siang et al. (2018) studied the effect of transitions between 4 cells (T1) and six cells (rosette) in the intercalation process. In the work of Wenzel et. al Wenzel et al. (2019), instead, structures composed of 4 and 5 cells are observed using a multiphase-field, active polar gel model.
@@
Here, for the first time, we make use of the Active Voronoi model to search for the minimal ingredients needed for the formation of rosette-like structures in 2D layers of cells. We isolate two key ingredients for observing such structures in monolayers through Voronoi models. The first ingredient is the breaking of the space translational invariance that generates topological defects. Those topological defects are candidates for the formation of rosette-like structures. However, without another dynamic ingredient that stabilizes those defects, rosettes are hindered by cell motility and deformability. We observe that self-alignment interactions, producing cell displacements in the same direction along which cells are elongated, are good candidates as minimal mechanisms for stabilizing those defects. We explored three paradigmatic cases in which the formation of rosettes is induced by breaking the spatial translational invariance of the system via the introduction of a privileged point: (i) the case of harmonic field, (ii) the case of a Lennard-Jones like potential that contains both, repulsion and attraction, and (iii) the case of a leader cell. In particular, harmonic and Lennard-Jones like potentials localized in one point of the space produce the formation of transient and lumen-centered rosettes, respectively. In the case of a leader cell, we considered a different intrinsic size with respect to the rest of the population to mimic the behavior of pinwheel rosettes Harding et al. (2014). For each of the three symmetry-breaking scenarios, we probed the effects of cell self-velocity, shape factor, alignment interaction, and persistence length, which are well-established control parameters in the cell tissue models, to isolate the minimal ingredients needed for observing stable rosette-like structures. We found that in the absence of nematic interaction (1) transient rosettes formation is permitted in confluent layers of deformable, slowly moving cells. The probability to observe rosette in motile, rigid cells scales with the intensity of the external potential/strength of the attractive force toward a specific point. (2) Alignment interaction both increases the probability of transient rosette formation and the size of lumen-centered ones. In particular, it allows cells to orderly dispose in multiple layers around the lumen. Finally, to test our model predictions, we performed experiments on iPSC-derived neural cells that are able to form rosettes after approximately eleven days of differentiation. Performing extensive fluorescent microscopy time lapses, we found (3) that neural rosettes are observed when the iPSCs populate the region of the parameter space our model predicts compatible for their formation.
@@
Figure 2: Trade-off between fluidity and stiffness a) From top to bottom, frequency of occurrence of cell as a function of their edges number, snapshot of representative stationary configurations, and trajectories upon varying the cell shape factor for a population of cells having self-propulsion velocity, $v0=0.5$ and persistence time, $\tau=50$ in presence of an external harmonic potential of intensity, $k=0.2$. b) Probability of observing rosette as a function of cell shape and cell motility for a population of cell subject to an external harmonic potential of intensity, $k=0.2$. Colors range from white to blue as the probability increases. c) Same as in panel b), but in presence of a harmonic potential of intensity, $k=0.3$.
@@
Figure 3: Alignment interaction favors rosette formation. a) Representative stationary configurations for populations of cells having different motility with ($J>0)$ and without ($J=0$) nematic interaction in presence of an external harmonic potential of intensity $k=0.2$. Observed rosettes are highlighted.b) Probability of observing rosette as a function of cell shape and nematic interaction strength for a population of cell subject to an external harmonic potential of intensity, $k=0.2$. Colors range from white to blue as the probability increases.
@@
Figure 4: Lumen centered rosettes. a) Snapshot of a stationary configuration obtained via the Active Voronoi model in presence of a Lennard-Jones potential. Cells forming the first, second and third layer around the central lumen are colored in green, gold and blue, respectively. b) Snapshots obtained for different values of the Lennard-Jones potential effective radius, $\sigma_{0}$. c) Number of cells in the first cellular layers as a function of time for different values of the cell target shape factor, $p_{0}$. d) Stationary number of cells in the first cellular layers as a function the cell target shape factor, $p_{0}$ and Lennard-Jones effective radius, $\sigma_{0}$ for two values of the Lennard-Jones potential strength, $\epsilon_{LJ}$. e) Same as in panel a) but with cells colored according to their elongation-alignment with respect to the lumen center. f) Snapshots obtained for different values of the cell target shape factor, $p_{0}$ in presence of a Lennard-Jones potential of effective radius, $\sigma_{0}=3.0$. g) Average polarization of cells in the first cellular layers as a function of time for different values of the cell target shape factor, $p_{0}$. h) Same as in panel d) but for the rosette polarization index. j) Same as in panel a) but for different values of the cell self-alignment. k) Boxplot representation of the distribution of the number of cells composing the rosette layers and their polarization indexes in absence and presence of self-alignment interaction.
@@
Figure 5: Leader cell-centered rosettes. a) Snapshots of stationary configurations obtained via the Active Voronoi model in presence of a leader cell of target area $A_{0}$. Top row displays snapshots for a population of cell having target shape factor $p_{0}=3.2$, while bottom row corresponds to cells with $p_{0}=3.8$. Cells forming the first, second and third layer around the central lumen are colored in green, gold and blue, respectively. b) Stationary number of cells in the first cellular layers (green shaded map) and polarization index (blue shaded map) as a function the cell target shape factor, $p_{0}$ and leader cell target area, $A_{0}$. c) Stationary polarization index of the first cellular layers as a function the cell target shape factor, $p_{0}$ and single cell velocity, $v_{0}$ for a population of cells in presence of a leader cell with target area, $A_{0}=10$. d) Boxplot representation of the distribution of the number of cells composing the rosette first layers in absence and presence of self-alignment interaction for different values of cells target shape factor, $p_{0}$. e) Same as in panel d) but for the polarization index of the cells in the rosette first layer.
@@ II.1 Rosette formation in the Voronoi Model
We consider a population of $N$ epithelial cells confined into a square box of side $L\!=\!\sqrt{N}$ with periodic boundary conditions. We indicate with ${\bm{r}}^{i}$ the position of the $i-$th cell center (with $i\!=\!1,..,N$) whose equation of motion reads
@@
Cell centers undergo Active Brownian dynamics in which the orientation of the self-propulsion $\theta_{i}$ follows a standard Langevin equation of the form $\dot{\theta}_{i}=\eta_{i}$, with $\langle\eta_{i}\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(s)\rangle=\frac{2}{\tau}\delta_{ij}\delta(t-s)$. The persistence time of the active motion, $\tau$ is the inverse of the rotational diffusivity coefficient $D_{r}$.
@@
$E_{v}^{i}\!=\!K_{p}(p_{i}-p_{0})^{2}+K_{A}(A_{i}-A_{0})^{2}$ (3)
In particular, $p_{i}$ and $A_{i}$ represent the perimeter and area of cell $i$, respectively, while $K_{p}$, $K_{a}$ are constants depending on the specific cell type. The area modulus $K_{a}$ determines how resistant a cell is to deviations from its preferred area. It penalizes deviations of the actual area of the cell from its target area $A_{0}$. The latter reflects the size that cells tend to maintain due to internal forces such as pressure, growth constraints, or packing density. Similarly, $K_{p}$ is the perimeter elasticity constant or perimeter modulus, which controls how resistant a cell is to changes in its perimeter with respect to its preferred perimeter $p_{0}$. Analogous to the preferred area $A_{0}$, the target perimeter reflects the cell’s preferred shape in terms of its perimeter and typically accounts for the contractility of the actomyosin network at the cell boundary. Without losing generality, one can set $K_{p}$, $K_{a}$, and $A_{0}$ to one, and keep $p_{0}$ as control parameters. In Figure 1a, we depict the relevant features of the model. A system subject only to vertex energy displays a phase transition between a solid and liquid-like behavior (see Figure 1b) for $p_{0}\sim 3.81$ (see for instance Bi et al. Bi et al. (2016)). To quantify the propensity of the model to form rosettes, we measured the frequency of occurrence of cells with less than 5 edges as proposed by Trichas et al. Trichas et al. (2012). As one can see from the bar plot in Figure 1b, the probability of observing rosettes is low as the system tends to organize in a mixture of pentagon and hexagon-like shapes, which minimize the Vertex energy in translational invariant space, i.e. flat ones. To induce rosette formation, we modified the topological features of the system in different manners, obtaining transient, lumen-centered, and pinwheel rosettes depending on the adopted defect source (see Figure 1i-iii).
@@ II.2 Formation of rosettes requires a tradeoff between tissue fluidity and single cell stiffness
Transient rosettes are observed in a wide range of cellular systems. To model their formation, we modified Eq. 2 by adding an external harmonic potential centered at a point C:
@@
The external potential may mimic the effect of the forces exerted by actin fibers, the attraction produced by the diffusion of specific molecules in the cellular micro-environment or/and the curvature of the organoid/embryo surface  Fernandez-Gonzalez and Zallen (2011). Figure 2 shows the results obtained simulating a population of 400 cells having self-propulsion velocity, $v0=0.5$ and persistence time, $\tau=50$ in presence of an external harmonic potential of intensity, $k=0.2$. Upon varying the cells target perimeter, i.e. their shape factor, the frequency of occurrence of cell having a number of edges lower than 5 increases with $p_{0}$ (see Figure 2a). Consequently, rosette like structures start appearing in the stationary configurations. Looking at the single cell trajectories, one clearly sees that cells in the rosette-forming region are highly motile, indicating that the formed rosettes are transient ones. Introducing an external field results in a fluidification of the cell layer even for cells with low target shape factor. To further investigate the effect of fluidity on rosette formation, we explored the behavior of the system upon changing single cell velocity, $v_{0}$. In absence of an external potential, increasing $v_{0}$ allows the tissue to undergo a solid to fluid transition Bi et al. (2016). Figure 2b displays the probability of observing cells having a number of edges lower than 5 as function of cell shape and cell motility for a population of cell subject to an external harmonic potential of intensity, $k=0.2$. While the probability systematically increases with cell shape deformability ($p_{0}$), motility dampens the capability of the cells to form rosettes. Indeed, varying the strength of the harmonic potential shows that there is a tradeoff between cell motility and harmonic strength that maximizes the probability of observing rosettes (see Figure 2c).
II.3 Nematic interactions boosts the probability of observing rosettes

### v2

@@ Abstract
Multicellular rosettes are ubiquitous structures in biological systems, observed in contexts ranging from morphogenesis to wound healing and cancer progression. While some molecular insights have been gained to explain the presence of these assemblies of cells around a common center, what are the tunable, global features that favor/hinder their formation is still largely unknown. Here, we made use of a Voronoi dynamical model to investigate the ingredients driving the emergence of rosettes characterized by different degrees of stability and organization. We found that (i) breaking the local spatial symmetry of the system, i.e., introducing curvature-inducing defects, allows for the formation of rosette-like structures (ii) whose probability of formation depends on the characteristics of the cellular layer. In particular, a trade-off between tissue fluidity and single-cell deformability dictates the assembly of transient rosettes, which are strongly stabilized in the presence of cell self-alignment interactions.
To test our model predictions, we performed fluorescence microscopy experiments on rosette-forming neural cell populations derived from induced pluripotent stem cells, finding significant agreement with the model predictions. Overall, our work may set the stage to gain an unifying understanding of the plethora of biophysical mechanisms involving the occurrence of rosette-like structures in physiology and their altered formation in pathology.
@@ I Introduction
Rosettes are polarized, transient epithelial structures, observed as intermediate states in several biological processes as different as morphogenesis, development, wound healing, and cancer Harding et al. (2014); Tetley et al. (2019); Gredler and Zallen (2023); Moran and Mosimann (2023). From a geometrical point of view, these structures can be identified in 2D sections of cellular layers as assemblies composed of five or more cells that interface at a central point Leng et al. (2021).
The mechanism through which such a configuration can be reached is proposed to be a topological rearrangement following either the disappearance of a cell (T2 transition) due to extrusion, ablation, or cell apoptosis Harding et al. (2014) or intercalation events where multiple T1 transitions lead to a 5 or more-fold vertex  Blankenship et al. (2006); Fletcher et al. (2014). The latter preserves the number of cells in the population.
Theoretical and numerical investigations so far have focused on studying the effect the presence of rosettes has on the tissue dynamics. Using modified vertex models, it has been shown that rosettes have an impact on tissue fluidity Yan and Bi (2019) and reorganization Trichas et al. (2012); Siang et al. (2018).
Indeed, transient rosettes are known to either resolve, producing tissue reorganizations Siang et al. (2018) or stabilize in lumen-centered polarized rosettes Harding et al. (2014), a critical process for organ development, including those of the kidneys, mammary glands, and neural tube Martin-Belmonte and Mostov (2008); Bryant and et al. (2010). Disregulation in such a process often results in severe developmental defects Avagliano et al. (2018).
Despite their centrality, the environmental conditions, i.e., the extracellular signals and cellular events, driving rosette formation are yet to be fully determined Miotto et al. (2023).
@@
Figure 1: Modeling rosette formation. a) Examples of stationary configurations in the solid and fluid-like regimes are shown with cells colored in blue and red, respectively. b) (top) Probability density distribution (pdf) of the cells’ edge lengths for the configurations shown in panel a). (bottom) Frequency of occurrence of cells as a function of their edge number. c) Sketches of three possible rosette conformations and the corresponding used model. (i) Transient rosette configurations, with 5 or more cells sharing a vertex, are simulated by breaking spatial symmetry via a harmonic potential acting on a given point. (ii) Rosettes with a central lumen are obtained via a Lennard-Jones potential energy originating in the center of the lumen, while (iii) pinwheel rosettes are simulated by inserting in the simulation a leader cell having a different target area.
Here, for the first time, we make use of the Active Voronoi model to search for the minimal ingredients needed for the formation of rosette-like structures in 2D layers of epithelial cells. In contrast to vertex models, where topological transitions have to be incorporated by hand, the Voronoi-based approach deals with transitions automatically as cell centers instead of cell vertices, making this coarse-grain model adapt to study rosettes dynamics without the need for parameter-richer models like those based on phase-fields Wenzel et al. (2019). Voronoi models are very effective as a coarse-grained description of confluent monolayers Park et al. (2015); Bi et al. (2016); Malinverno et al. (2017); Giavazzi et al. (2018); Li and Ciamarra (2018). However, additional ingredients are required for triggering rosette formation in Voronoi models.
@@
We isolate two key ingredients for observing such structures in monolayers through Voronoi models. The first ingredient is the breaking of the space translational invariance that generates topological defects. Those topological defects are candidates for the formation of rosette-like structures. However, without another dynamic ingredient that stabilizes those defects, rosettes are hindered by cell motility and deformability. We observe that self-alignment interactions, producing cell displacements in the same direction along which cells are elongated, are good candidates as minimal mechanisms for stabilizing those defects. We explored three paradigmatic cases in which the formation of rosettes is induced by breaking the spatial translational invariance of the system via the introduction of a privileged point: (i) the case of harmonic field, (ii) the case of a Lennard-Jones (LJ) like potential that contains both, repulsion and attraction, and (iii) the case of a leader cell. In particular, harmonic and LJ-like potentials localized at one point of the space produce the formation of transient and lumen-centered rosettes, respectively. In the case of a leader cell, we considered a different intrinsic size with respect to the rest of the population, mimicking the behavior of pinwheel rosettes Harding et al. (2014). For each of the three symmetry-breaking scenarios, we probed the effects of cell self-velocity, shape index, self-alignment interaction, and persistence length, which are well-established control parameters in cell tissue models, to isolate the minimal ingredients needed for observing stable rosette-like structures. We found that in the absence of self-alignment interaction, (1) transient rosette formation is permitted in confluent layers of deformable, slowly moving cells. The probability of observing rosettes in motile, rigid cells scales with the intensity of the external potential/strength of the attractive force toward a specific point. (2) Alignment interaction both increases the probability of transient rosette formation and the size of lumen-centered ones. In particular, it allows cells to be orderly disposed in multiple layers around the lumen. Finally, to test our model predictions, we performed experiments on iPSC-derived neural cells that are able to form rosettes after approximately eleven days of differentiation. Performing extensive fluorescent microscopy time lapses, we found (3) that neural rosettes are observed when the iPSCs populate the region of the parameter space our model predicts to be compatible for their formation.
@@
Figure 2: Trade-off between fluidity and stiffness a) From top to bottom, frequency of occurrence of cell as a function of their edges number, snapshot of representative stationary configurations, and trajectories upon varying the cell shape factor for a population of cells having self-propulsion velocity, $v0=0.5$ and persistence time, $\tau=50$ in presence of an external harmonic potential of intensity, $k=0.2$. b) Probability of observing rosette as a function of cell shape and cell motility for a population of cells subject to an external harmonic potential of intensity, $k=0.2$. Colors range from white to blue as the probability increases. c) Same as in panel b), but in the presence of a harmonic potential of intensity, $k=0.3$.
@@
Figure 3: Alignment interaction favors rosette formation. a) Representative stationary configurations for populations of cells having different motility with ($J>0)$ and without ($J=0$) self-alignment interaction in the presence of an external harmonic potential of intensity $k=0.2$. Observed rosettes are highlighted.b) Probability of observing rosette as a function of cell shape and self-alignment interaction strength for a population of cells subject to an external harmonic potential of intensity, $k=0.2$. Colors range from white to blue as the probability increases.
@@
Figure 4: Effect of model parameters on lumen-centered rosettes. a) Snapshot of a stationary configuration obtained via the Active Voronoi model in the presence of a Lennard-Jones potential. Cells forming the first, second, and third layers around the central lumen are colored in green, gold, and blue, respectively. b) Snapshots obtained for different values of the LJ potential effective radius, $\sigma_{0}$. c) Number of cells in the first cellular layers as a function of time for different values of the cell target shape factor, $p_{0}$. d) Stationary number of cells in the first cellular layers as a function of the cell target shape factor, $p_{0}$, and LJ effective radius, $\sigma_{0}$, for two values of the Lennard-Jones potential strength, $\epsilon_{LJ}$. e) Same as in panel a) but with cells colored according to their elongation-alignment with respect to the lumen center. f) Snapshots obtained for different values of the cell target shape factor, $p_{0}$, in the presence of a Lennard-Jones potential of effective radius, $\sigma_{0}=3.0$. g) Average polarization of cells in the first cellular layers as a function of time for different values of the cell target shape factor, $p_{0}$. h) Same as in panel d) but for the rosette polarization index. j) Same as in panel a) but for different values of the cell self-alignment. k) Boxplot representation of the distribution of the number of cells composing the rosette layers and their polarization indexes in the absence and presence of self-alignment interaction.
@@
Figure 5: Comparison between lumen and leader cell-centered rosettes. a) Snapshots of stationary configurations obtained via the Active Voronoi model in the presence of a leader cell of target area $A_{0}$. The top row displays snapshots for cells having a target shape factor $p_{0}=3.2$, while the bottom row corresponds to cells with $p_{0}=3.8$. Cells forming the first, second, and third layers around the central lumen are colored in green, gold, and blue, respectively. b) Stationary number of cells in the first cellular layers (green shaded map) and polarization index (blue shaded map) as a function of the cell target shape factor, $p_{0}$, and leader cell target area, $A_{0}$. c) Stationary polarization index of the first cellular layers as a function of the cell target shape factor, $p_{0}$, and single cell velocity, $v_{0}$, for a population of cells in the presence of a leader cell with target area, $A_{0}=10$. d) Boxplot representation of the distribution of the number of cells composing the rosette first layers in the absence and presence of self-alignment interaction for different values of the cells’ target shape factor, $p_{0}$. e) Same as in panel d) but for the polarization index of the cells in the rosette first layer.
@@ II.1 Rosette formation in the Voronoi Model
We consider a population of $N$ epithelial cells confined in a square box of side $L\!=\!\sqrt{N}$ with periodic boundary conditions. We indicate with ${\bm{r}}^{i}$ the position of the $i$-th cell center (with i=1,..,Ni\!=\!1,..,N) whose equation of motion reads Bi et al. (2016)
@@
Cell centers undergo Active Brownian dynamics in which the orientation of the self-propulsion $\theta_{i}$ follows a standard Langevin equation of the form $\dot{\theta}_{i}=\eta_{i}$, with $\langle\eta_{i}\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(s)\rangle=\frac{2}{\tau}\delta_{ij}\delta(t-s)$. The persistence time of the active motion, $\tau$, is the inverse of the rotational diffusivity coefficient $D_{r}$.
@@
$E_{v}^{i}\!=\!K_{p}(p_{i}-p_{0})^{2}+K_{A}(A_{i}-A_{0})^{2}\;.$ (3)
In particular, $p_{i}$ and $A_{i}$ represent the perimeter and area of cell $i$, respectively, while $K_{p}$ and $K_{a}$ are constants depending on the specific cell type. The area modulus $K_{a}$ determines how resistant a cell is to deviations from its preferred area. It penalizes deviations of the actual area of the cell from its target area $A_{0}$. The latter reflects the size that cells tend to maintain due to internal forces such as pressure, growth constraints, or packing density. Similarly, $K_{p}$ is the perimeter elasticity constant or perimeter modulus, which controls how resistant a cell is to changes in its perimeter with respect to its preferred perimeter $p_{0}$. Analogous to the preferred area $A_{0}$, the target perimeter reflects the cell’s preferred shape in terms of its perimeter and typically accounts for the contractility of the actomyosin network at the cell boundary. Without losing generality, one can set $K_{p}$, $K_{a}$, and $A_{0}$ to one, and keep $p_{0}$ as a control parameter. In Figure 1a, we depict the relevant features of the model. A system subject only to vertex energy displays a phase transition between a solid and liquid-like behavior (see Figure 1b) for $p_{0}\sim 3.81$ (for details, see  Bi et al. (2016)). To quantify the propensity of the model to form rosettes, we measured the frequency of occurrence of cells with less than 5 edges, $P_{r}$, along the line of Trichas et al. Trichas et al. (2012). As one can see from the bar plot in Figure 1b, the probability of observing rosettes is low as the system tends to organize in a mixture of pentagon and hexagon-like shapes, which minimize the Vertex energy in translational invariant space, i.e., flat ones. To induce rosette formation, we modified the topological features of the system in different manners, obtaining transient, lumen-centered, and pinwheel rosettes depending on the adopted defect source (see Figure 1i-iii).
@@ II.2 Formation of rosettes requires a tradeoff between tissue fluidity and single cell stiffness
To promote the formation of transient rosettes, we modified Eq. 2 by adding an external harmonic potential centered at a point C:
@@
The external potential may mimic the effect of the forces exerted by actin fibers, the attraction produced by the diffusion of specific molecules in the cellular micro-environment, and/or the curvature of the organoid/embryo surface  Fernandez-Gonzalez and Zallen (2011). Figure 2 shows the results obtained simulating a population of 400 cells having self-propulsion velocity, $v0=0.5$, and persistence time, $\tau=50$, in the presence of an external harmonic potential of intensity, $k=0.2$. Upon varying the cells’ target perimeter, i.e., their shape index, the frequency of occurrence of cells having a number of edges lower than 5, $P_{r}$, increases with $p_{0}$ (see Figure 2a). Consequently, rosette-like structures start appearing in the stationary configurations. Looking at the single cell trajectories, one clearly sees that cells in the rosette-forming region are highly motile, indicating that the formed assemblies are transient ones. Note that introducing an external field, i.e., a curvature, results in a fluidification of the cell layer even for cells with low target shape factor Sussman (2020); De Marzio et al. (2025). To further investigate the effect of fluidity on rosette formation, we explored the behavior of the system upon changing cell self-propulsion velocity, $v_{0}$. In the absence of an external potential, increasing $v_{0}$ allows the tissue to undergo a solid to fluid transition Bi et al. (2016). Figure 2b displays $P_{r}$ as a function of cell shape and cell motility for a population of cells subject to an external harmonic potential of intensity, $k=0.2$. While the probability systematically increases with cell shape deformability ($p_{0}$), motility dampens the capability of the cells to form rosettes. Indeed, varying the strength of the harmonic potential shows that there is a tradeoff between cell motility and harmonic strength that maximizes the probability of observing rosettes (see Figure 2c).
II.3 Self-alignment interactions boost the probability of observing rosettes